# IoT Ground Sensing Systems for Early Wildfire Detection: Technologies, Challenges and Opportunities

### v1

@@ V Potential Framework and Possible Extension
Despite the challenges encountered, the use of an IoT sensing system can be a valuable tool for early wildfire detection. By integrating detection using environmental data with other imaging detection platforms, such as satellites, UAVs, watchtowers and vision-based IoT, precise early warning systems can be established. In this section, our aim is to present a potential framework for the development of IoT sensing systems addressing issues we discussed in the previous section, with a particular emphasis on reducing energy consumption and dealing with the trade-off when selecting sensors and detection algorithms, through the use of multi-criteria decision analysis (MCDA). The potential for further development and exploration of the IoT ground sensing system for the early detection of wildfires is also discussed.
V-A Potential Framework
@@
To reserve energy in the IoT sensing system, it is suggested to use energy harvesting techniques, such as solar energy harvesting, to recharge the battery of EDs. The battery should last for up to the next recharge period. During periods characterised by clear skies, the energy harvested from the solar panel is likely to fully charge the battery. As mentioned in Section II, MoS and pellistor-type sensors require heating, which can lead to a shorter lifespan of EDs and higher maintenance costs. Therefore, alternative sensing technologies, such as NDIR and FET, which consume only a few microwatts per second, should be taken into account. To further reduce energy consumption in IoT EDs, the sensors can be set to sample at different rates according to weather conditions or the calculated moisture content of fire fuel related to the higher risk of fire on a given day. For example, during extreme heat and aridity conditions, it is advisable to increase the sampling rate to collect a larger number of readings from the devices while reducing it under opposite conditions to conserve the battery. This approach has already been implemented in various projects, including [80, 198].
@@
In terms of the deployment height and orientation of IoT EDs, based on our experiments and insights from Dryad study [29] suggest that end devices should be mounted on trees at a height of 2-3 metres to optimise the detection of TVOC and CO in the event of a wildfire. In addition, elevated placement of EDs also improves the LoS between EDs and the gateway. Our experiments have shown that MoS sensors can detect a small fire burning in a 0.7 square metre fire pit within a few minutes from a distance of 50 metres when placed downwind. This implies that the inter-node distances for a grid topology of EDs can be metres. The unpredictability of wildfire ignition points in forested areas makes it difficult to determine the best orientation for individual sensors to maximise their detection capabilities. These sensors are similar to directional antennas, capable of sensing only in one direction. Our experiments have shown that when the sensor is not pointed towards the fire source, the detected gas concentration can be reduced to only 10 of its original value when facing the ignition point. To address this issue, [91] proposed a method that involves placing the sensors in a partially enclosed cage, with the opening facing the ground. This design effectively turns a unidirectional sensor into an omnidirectional one. Our tests confirm that this configuration enables the detection of approximately 50 of the gas concentration that would be perceived at maximum reception when the sensor is aligned with the fire source.
@@
To address the challenge of packet collision and the energy consumption that follows from packet loss, it is recommended to reduce the packet size and transmission frequency. Instead of sending all sensor information back to the server for investigation, research has shown that statistical threshold detection and basic ML algorithms such as KNN or linear SVM can be used for preliminary filtering [169, 170, 89, 181, 96]. When EDs detect a series of alerts related to meteorological and gas parameters, they should trigger an alarm and notify fire departments. Then, EDs can send their observational data to a central server, where more advanced DL techniques, such as RNN or LSTM networks, can be used for thorough cross-verification. The central server can apply spatial correlation strategies to accurately identify the possibility of a fire hazard [186]. After confirming a fire incident, camera modules on site can be used to acquire visual evidence for further examination. In addition, direct to satellite communication is viable for remote, network-lacking areas. Although it increases gateway energy consumed, costs can be reduced compared to packet relaying. In order to enhance the dependability of the system and reduce false positives, the generated alert can also be confirmed by cross-validating with watchtowers, UAVs, or satellite entities. In conclusion, the integration and centralisation of IoT technology, UAVs, watchtowers and satellite systems should be coordinated and monitored by fire control centres.
@@
In this section, we have mentioned a range of strategies to enhance the performance of an IoT ground sensing system, focusing on prolonging the lifespan of EDs, optimising detection capabilities and minimising transmission delays. Despite progress in the IoT sensing system, there are still fundamental restrictions due to hardware performance limitations in EDs and energy consumption. This ongoing challenge requires balancing the trade-off between energy consumption, detection accuracy and detection latency. Therefore, the MCDA framework can be used to balance the specified requirements. In a basic MCDA model, assigning a weight to each evaluation criterion enables the calculation of the overall effectiveness, as the weighted sum of all criteria. In this case, the contribution of each criterion is directly proportional to its assigned weight. Additionally, some criteria, such as the latency in detection or the sensor’s recovery time, can be given an exponential weight, thus introducing a combination of linear and exponential weights into the analysis. This modified MCDA cost can be expressed by the following equation,
@@
where the values of the criteria represents the criteria contributing linearly to and represents the criteria contributing exponentially to the cost. and are the corresponding linear weights and exponential weights, respectively. The indices and range from 1 to and 1 to respectively, with being the total number of evaluation criteria. This MCDA framework can be customised to assist in the selection of a suitable sensor based on factors such as cost, energy consumption, response recovery time and operational environmental ranges. Additionally, it can be used to determine the most effective anomaly detection method, whether point-wise, time-series, or ML/DL detection techniques. Furthermore, unsupervised detection algorithms cannot be evaluated using the F1 scoring system as in supervised learning, so the MCDA can also be used to effectively compare the two methods. This can be done by taking into account criteria such as the delay from the ignition point until detection and the number of false alarms. Last but not least, the framework can be used to evaluate the effectiveness of on-board processing in EDs or processing on the centralised server.
V-B Possible Extensions
The effective detection of wildfires is based on the use of multiple sensors, each capable of detecting different fire indicators such as heat, TVOC and CO. Integration of data from various sensors allows for a more accurate assessment of the fire situation and facilitates informed decision-making on response strategies. In a study by [199], the authors summarise the correlation analysis between various types of sensor data, revealing that TVOC, CO, and PM exhibit limited correlation with fire alarms. Consequently, more research effort is needed to determine the optimal combination of sensor inputs that can effectively enhance the accuracy of ML/DL detection algorithms. Furthermore, a comprehensive examination of the sensor system’s topological configuration is also important, particularly in the context of IoT-based wildfire detection algorithms deployed in forested areas where the fire source obstructs NLoS between most sensors. Understanding whether the behaviour of the system differs from that of open outdoor areas is crucial, including aspects such as optimal sensor height and topology. Lastly, performing an analysis of the energy consumption of the solar battery system and exploring novel techniques to reach the maximum power point of the charging circuitry can significantly enhance the performance of energy charging. These areas of investigation are essential to improve the overall effectiveness and sustainability of the wildfire detection system.
@@ VI Conclusion
An overview of the state-of-the-art of IoT ground sensing wildfire detection systems is presented in this article. The underlying principles of various vision-based detection algorithms, as well as the working principles of multiple environmental sensing units and various detection algorithms employed in ground sensing anomaly detection, are explained. A detailed discussion of the challenges encountered during the development of an IoT ground sensing system for wildfire detection is also provided. Based on the findings of the literature review and our own experimental observations, a potential framework for the development of a robust and reliable detection system is proposed, along with suggestions for future research possibilities.

### v2

@@ V Suggestions for Future Implementation
Despite the challenges encountered, the use of IoT sensing systems can be a valuable tool for early wildfire detection. By integrating detection using environmental data with other imaging detection platforms, such as satellites, UAVs, watchtowers and vision-based IoT, precise early warning systems can be established. In this section, our aim is to provide recommendations for the development of IoT sensing systems addressing issues we discussed in the previous section, with particular emphasis on reducing energy consumption and dealing with the trade-off when selecting sensors and detection algorithms, through the use of a decision-making model. The potential for further development and exploration of IoT ground sensing systems for the early detection of wildfires is also discussed. Fig. 7 illustrates the trade-off relationship or the so-called ’dilemma’ that arises when implementing the IoT ground sensing system for early wildfire detection. The critical factors to consider for IoT engineers deploying such a system are energy consumption, system accuracy, and detection delay. Suggestions are provided to enhance the capacity of the IoT system in one aspect without significantly compromising other aspects. These suggestions are based on a critical review of the literature combined with the preliminary result of our ongoing research.
V-A Addressing Energy Consumption
To reserve energy in the IoT sensing system, it is suggested to use energy harvesting techniques, such as solar energy harvesting, to recharge the ED battery. The battery should last up to the next recharge period. During periods characterised by clear skies, the energy harvested from the solar panel is likely to fully charge the battery. Among all energy harvesting methods, solar energy harvesting stands out as the most practical choice for powering remote IoT systems because of its well-established nature in supplying remote power and its reliability in forested areas, as well as the effective deployment of solar hardware. To investigate the effectiveness of solar power harvesting, we measured the daily energy consumption of the end node and the gateway, selected appropriate solar panel controllers and photovoltaic panels (10 watts for the end node, 30 watts for the gateway) based on an assumed 80% exposure to the mean daily irradiance, and verified daily energy recharge of the system. Fig. 8 shows the accumulation of energy by the photovoltaic modules of an ED and the battery capacity over a couple of days, thus demonstrating the practicality of incorporating the solar energy harvesting system within the ED framework. To further increase the efficiency of solar energy harvesting, different research projects have suggested and used maximum power point tracking (MPPT) techniques [210, 211]. The MPPT module can auto-track the maximum point of output voltage to optimize the energy conversion. Typically, MPPT can improve energy conversion efficiency by about 10 to 30 compared to systems without MPPT under various conditions, in non-ideal circumstances such as partial shading, cloudy weather, or varying temperatures. In a recent study [212], the radiative cooling film is combined with a solar panel to further increase the efficiency of energy harvesting by adding an energy harvesting feature through radiative cooling to a solar harvesting system and has been proven to outperform the traditional method. However, further investigations are imperative to determine the maximum charging range and to evaluate the efficacy of such applications in outdoor environments.
@@ V-A Addressing Energy Consumption
Beyond energy harvesting, as highlighted in Section II, both MoS and pellistor-type sensors require heating, potentially shortening the lifespan of EDs and increasing maintenance expenses. Therefore, alternative sensing technologies, such as NDIR and FET, which consume only a few microwatts per second, should be taken into account. To further reduce energy consumption in IoT EDs, the sensors can be set to sample at different rates according to weather conditions or the calculated moisture content of fire fuel related to the higher risk of fire on a given day. For example, during extreme heat and aridity conditions, it is advisable to increase the sampling rate to collect a larger number of readings from the devices while reducing it under opposite conditions to conserve the battery. This approach has started to be adopted in recent works such as [81].
V-B Improving Sensing Accuracy
@@
While the effectiveness of the IoT wildfire detection system is highly dependent on the complexity of its detection algorithm, leading to a decision on the selection of detection algorithms, the accuracy is also affected by the deployment approaches used for the sensor nodes. By strategically placing sensors at specific heights and aligning them correctly, it is possible to optimise the detection range, leading to increased sensitivity and ultimately reducing the rate of missed detections. In terms of height and orientation of IoT ED deployment, based on our experiments and the insights from Dryad study [37] suggest that the end devices should be mounted on trees at a height of 2-3 metres to optimise the detection of TVOC and CO2 in the event of a wildfire. Furthermore, the elevated placement of EDs also improves the LoS between EDs and the gateway. Our experiments have shown that MoS sensors can detect a small fire burning in a 0.7 square metre fire pit within a few minutes from a distance of 50 metres when placed downwind. This implies that the inter-node distances for a grid topology of EDs can be metres. The unpredictability of wildfire ignition points in forested areas makes it difficult to determine the best orientation for individual sensors to maximise their detection capabilities. These sensors are similar to directional antennas, capable of sensing only in one direction. Our experiments have shown that when the sensor is not pointed towards the fire source, the detected gas concentration can be reduced to only 10 of its original value when facing the ignition point. To address this issue, [95] proposed a method that involves placing the sensors in a partially enclosed cage, with the opening facing the ground. This design effectively turns a unidirectional sensor into an omnidirectional one. Our tests confirm that this configuration enables the detection of approximately 50 of the gas concentration that would be perceived at maximum reception when the sensor is aligned with the fire source. Nevertheless, analysis of the system topology is also crucial. Specifically, adjustments may be necessary when implementing an IoT system in certain forested regions where the presence of fire obstructs the LoS between sensors and it is critical to determine if the system’s performance varies from that in open outdoor spaces, taking into account factors like the best sensor elevation and configuration. The effective detection of wildfires is based on the use of multiple sensors, each capable of detecting different fire indicators such as heat, TVOC and CO2. Additionally, to improve the accuracy of detection systems, an effective approach involves incorporating various detection dimensions. This includes expanding the types of input data for decision-making, such as including temperature, RH, and air pressure in the detection model, given that certain activities by humans or animals can cause an increase in RH levels, but in a study by [213], the authors summarise the correlation analysis between various types of sensor data, revealing that TVOC, CO2, and PM exhibit limited correlation with fire alarms. Consequently, more research effort is needed to determine the optimal combination of sensor inputs that can effectively enhance the accuracy of ML/DL detection algorithms.
V-C Reducing Detection Delay
@@
The strategies recommended in the preceding section not only enhance detection accuracy but also aid in minimising detection latency by improving sensor sensitivity. However, further measures are needed to mitigate delays resulting from packet loss between EDs and the gateway, and subsequently to the cloud server. To address the challenge of detection delay and energy consumption resulting from packet loss, it is recommended to reduce packet size and transmission frequency. Instead of sending all sensors data back to the server for investigation, research has shown that statistical threshold detection and basic ML algorithms such as KNN or linear SVM can be used for preliminary filtering [174, 175, 93, 186, 100]. For cloud-based detection, when EDs identify a sequence of alerts related to meteorological and gas parameters, they should activate an alarm and inform fire departments. Subsequently, EDs will transmit their observational data to a central server, where more sophisticated DL techniques, such as RNN or LSTM networks, can be employed for comprehensive cross-verification. The central server can apply spatial correlation strategies to accurately identify the possibility of a fire hazard [192]. Once the fire incident is confirmed, camera modules on site can be used to acquire visual evidence for further examination. In addition, direct satellite communication is viable for remote areas that lack network connectivity. Although it increases gateway energy consumed, costs can be reduced compared to packet relaying. To enhance the dependability of the system and reduce false positives, the generated alert can also be confirmed by cross-validating with watchtowers, UAVs, or satellite entities. In conclusion, the integration and centralisation of IoT technology, UAVs, watchtowers and satellite systems should be coordinated and monitored by fire control centres.
V-D Trade-off and Decision Making
@@
In this section, we have mentioned a range of strategies to enhance the performance of an IoT ground sensing system, focusing on prolonging the lifespan of EDs, optimising detection capabilities and minimising transmission delays. Despite progress in the IoT sensing system, there are still fundamental restrictions due to hardware performance limitations in EDs and energy consumption. This ongoing challenge requires balancing the trade-off between energy consumption, detection accuracy and detection latency. Hence, creating a metric for evaluating these aspects and determining the appropriateness levels can assist IoT engineers in making well-informed decisions. To this end, the multi-criteria decision analysis (MCDA) model is an appropriate decision making framework that can be used to weigh the defined criteria. In a basic MCDA model, assigning a weight to each evaluation criterion enables the calculation of the overall effectiveness, as the weighted sum of all criteria. In this case, the contribution of each criterion is directly proportional to its assigned weight. Additionally, some criteria, such as detection latency or sensor recovery time, can be given an exponential weight, thus introducing a combination of linear and exponential weights into the analysis. This modified MCDA cost can be expressed by the following equation,
@@
where the values of the criteria represents the criteria contributing linearly to and represents the criteria contributing exponentially to . and are the corresponding linear and exponential weights, respectively. The indices and range from 1 to and 1 to respectively, with being the total number of evaluation criteria. This MCDA framework can be customised to assist in the selection of a suitable sensor based on factors such as cost, energy consumption, response time, and operational environmental ranges.
Table VII presents an illustration of how MCDA can be used to evaluate and compare different gas/smoke sensors. In this particular example, we assume that energy consumption, sensor cost, and sensor size have a linear weighted impact on overall cost. Specifically, we assume that an increase of 50 mW in power consumption, an additional $20 in sensor cost, and an increase of 5 in sensor size are deemed to cause a similar level of challenge in hardware implementation or design. Consequently, these factors are assigned costs of 0.02, 0.05, and 0.2, respectively. As minimizing detection delays is of utmost importance, we assign it an exponential weighting. We consider the sensor to be satisfactory if it can detect smoke as changes in CO2, TVOC, or particulates in 90 seconds, with exponential weight equal to . The SGP30 sensor has the lowest weighted sum cost among the sensors, making it the most suitable candidate for this particular MCDA.
Additionally, the MCDA can be used to determine the most effective anomaly detection method, whether point-wise, time-series, or ML/DL detection techniques. As unsupervised detection algorithms cannot be evaluated using the F1 scoring system as in supervised learning, the MCDA can also be used to effectively compare the two methods. This can be done by taking into account criteria such as the delay from the ignition point until detection and the number of false alarms. Last but not least, the framework can be used to evaluate the effectiveness of on-board processing in EDs or processing on the centralised server.
@@ VI Conclusion
An overview of the state-of-the-art of IoT ground sensing wildfire detection systems is presented in this article. The underlying principles of various vision-based detection algorithms, as well as the working principles of multiple environmental sensing units and various detection algorithms employed in ground sensing anomaly detection, are explained. A detailed discussion of the challenges encountered during the development of an IoT ground sensing system for wildfire detection is also provided. Based on the findings of the literature review and our own experimental observations, suggestions for the development of a robust and reliable IoT early wildfire detection system are proposed, along with suggestions for future research possibilities.